# Conformal Online Learning of Deep Koopman Linear Embeddings

### v1

@@ 1 Introduction
Understanding the evolution of complex systems over time is essential in numerous disciplines, including robotics (Bruder2021), finance (mann2016dynamic), physics (kaptanoglu2020characterizing), chemistry (klus2020kernel) and biology (Hasnain2020). A particularly powerful framework for this analysis is provided by operator-theoretic approaches, which recast nonlinear dynamics into linear evolution in function space. Among these, the Koopman operator (budivsic2012applied; mauroy2020koopman) plays a central role: it describes the progression of measurement functions (or observables) defined over the state space, yielding an infinite-dimensional linear representation of nonlinear systems. Its spectral decomposition offers a principled way to characterize the long-term behavior and underlying structure of the dynamics (Bruton2022ModernKoopman).
@@
Although the Koopman operator is infinite-dimensional, several numerical methods have been developed to approximate it in a finite-dimensional setting. Notable among these are Dynamic Mode Decomposition (DMD) (Rowley2009DMD) and its nonlinear extensions, such as Extended DMD (EDMD) (Williams2015EDMD) and its variants (see (Jin2024) and references therein). Recent work has also proposed truncated Signature features rooted in path theory (chretien2025sigkoop). Another line of research consists of kernel learning formulations in reproducing kernel Hilbert spaces (kostic2022rkhs; pmlr-v202-hou23c). In addition, deep learning techniques have been leveraged to learn expressive Koopman-invariant representations, using neural networks and autoencoder architectures to estimate significant observables (Li2017Dictionary; wehmeyer2018time; Lusch2018Universal; yeung2019learning; Otto2019Autoencoder). For example, xu2025reskoopnet addresses this by explicitly minimizing the spectral residual. The stochastic Koopman operator extends the classical Koopman framework to random dynamical systems, capturing the evolution of observables under stochastic influences (vcrnjaric2020SKO; kostic2023sharp; xu2025sdmd).
@@
Most existing methods, however, operate in the offline setting, assuming access to the entire dataset in advance. Yet, in many applications—such as online monitoring, adaptive control, or real-time forecasting—data arrive sequentially, and the system may evolve in a non-stationary fashion (Korda2018). Several online methods have been proposed for this task. For instance, Online DMD (Zhang2018OnlineDMD) and Online EDMD (sinha2019rRecursiveEDMD; sinha2023online) incrementally update Koopman operator estimates as new data arrives. However, these methods typically rely on linear observables or fixed dictionaries, limiting their expressiveness. More recent methods using neural networks (e.g. (liang2022online; Hao2024DKLT)) lack principled learning strategies and often rely on retraining the model using a fixed number of steps, regardless of whether the update is necessary. These limitations highlight the need for online learning strategies that are not only memory-efficient but also adaptive in order to update models only when required by the incoming data.
@@
We address this need by introducing Conformal Online Learning of Koopman embeddings (COLoKe), whose high-level principle is sketched in Figure 1. Our approach combines deep Koopman representation learning with a novel repurposing of conformal prediction principles to decide when to adapt the model as data arrive in a streaming fashion. Updates are triggered only when necessary, reducing both computational burden and overfitting. This contributes to the growing body of work on online Koopman learning (sinha2023online; Loya2024Online; Mazouchi2023BatchOnline). Our method differs by enabling adaptive embeddings, built-in reconstruction, and real-time updates while remaining memory-efficient (see Table 1). To the best of our knowledge, COLoKe is the first principled approach for online learning driven by conformal-based updates.
@@
No need
for history
Online
update
Adaptative
embedding
Built-in
reconstruction
ODMD (Zhang2018OnlineDMD)
R-EDMD (sinha2019rRecursiveEDMD; sinha2023online)
DKLT (Hao2024DKLT) batch only
BatchOnline(Mazouchi2023BatchOnline) batch only
@@ 2.1 Data-driven learning of the Koopman operator
which describes the evolution of observables along trajectories of the system. Importantly, $\mathcal{K}$ is a linear operator, even when $T$ is nonlinear, making it a powerful tool for the spectral analysis of nonlinear dynamics (Bruton2022ModernKoopman). In particular, if $\varphi\in\mathcal{F}$ is an eigenfunction of $\mathcal{K}$ with eigenvalue $\lambda\in\mathbb{C}$, i.e., $\mathcal{K}\varphi=\lambda\varphi$, then along a trajectory $(x_{t})$, the observable evolves linearly: $\varphi(x_{t})=\lambda^{t}\varphi(x_{0})$. It follows that, when a set of eigenfunctions $\{\varphi_{1},\dots,\varphi_{L}\}$ defines an injective embedding of the state space, the system can be linearized via the coordinate transformation $x\mapsto\left(\varphi_{1}(x),\dots,\varphi_{L}(x)\right)$. In this lifted space, the nonlinear dynamics evolve linearly, providing a compelling framework for Koopman-based analysis and control (Korda2018; mauroy2020koopman).
@@
where $\Phi(x)=[f_{1}(x),\dots,f_{m}(x)]^{\top}$ denotes the lifted representation of the state. This insight motivates the search for low-dimensional Koopman-invariant subspaces that admit such linear representations. Methods like EDMD (Williams2015EDMD) approximate this setting by fixing a dictionary $\{f_{i}\}_{i=1}^{m}$, but their performance is limited by the expressiveness and suitability of the chosen observables. To overcome this limitation, recent approaches (takeishi2017learning; Lusch2018Universal; yeung2019learning; Otto2019Autoencoder) propose to learn both the feature map $\Phi$ and the linear operator $K$ jointly using neural networks, leading to deep Koopman embeddings.
@@ 2.2 Conformal prediction for online data
Conformal prediction (vovk1999machine; vovk2005algorithmic; romano2019conformalized; angelopoulos2021gentle) is a distribution-free framework for uncertainty quantification that constructs valid prediction sets with finite-sample guarantees.
Given past input–output pairs $\{(x_{i},y_{i})\}_{i=1}^{t-1}$, a model produces a prediction $\hat{y}_{t}$ for a new input $x_{t}$, and assigns a conformity score $s(x_{t},y)$ to each candidate output $y$. The conformal prediction set is defined as $C_{t}=\left\{y\in\mathcal{Y}\;\middle|\;s(x_{t},y)\leq q_{t}\right\}$, where $q_{t}$ is a quantile calibrated to ensure $\mathbb{P}(y_{t}\in C_{t})\geq 1-\alpha$. The set $C_{t}$ is thus interpreted as a set of plausible outputs: it contains all candidate values of $y$ that are deemed sufficiently "conformal" (i.e., not too surprising) with respect to the current model and past observations. While this framework is powerful and requires no distributional assumptions beyond exchangeability, it is not directly applicable to time series or online learning, where data are typically non-exchangeable. In such settings, standard conformal methods may yield miscalibrated or overly conservative intervals. Addressing this challenge has motivated the development of adaptive conformal approaches that can track distribution shifts over time (Gibbs2021; xu2021conformal; Angelopoulos2023ConformalPID). In particular, Conformal PID Control was recently introduced in Angelopoulos2023ConformalPID as a dynamically calibrated version of conformal prediction. Here, “PID” refers to the use of Proportional, Integral, and (optionally) Derivative feedback terms—standard components in control theory—used to adaptively adjust the prediction threshold. Rather than fixing the quantile threshold $q_{t}$ in advance, the method updates it online in response to conformity violations. Among its variants, we consider the conformal PI control scheme: after observing whether $y_{t}\in C_{t}$, a binary error signal $e_{t}=\mathbf{1}\{y_{t}\notin C_{t}\}$ is computed, and the threshold is adjusted via:
@@
where $\gamma>0$ is a learning rate and $r_{t}$ is a nonlinear saturation function (Angelopoulos2023ConformalPID) acting as an integral correction term. This PI-style control loop allows the prediction region to adaptively expand or contract to maintain the desired coverage. Under mild regularity conditions, it ensures that the empirical coverage converges, i.e., $\frac{1}{T}\sum_{t=1}^{T}e_{t}\to\alpha$ as $T\to\infty$.
@@ 3.1 From online learning…
At each time step $t$, we denote by $\Phi_{\theta_{t}}:\mathcal{X}\to\mathbb{C}^{m}$ the current feature map and by $K_{t}\in\mathbb{C}^{m\times m}$ the corresponding finite-dimensional Koopman operator. Following (Li2017Dictionary), we also advocate to enforce interpretability and preserve part of the original state, by designing the feature map to include both the identity and a learnable nonlinear component, i.e.,
@@
This formulation is motivated by two design principles. First, as shown by Otto2019Autoencoder, multi-step prediction promotes the identification of persistent spectral modes and approximate Koopman eigenfunctions, thereby improving long-term prediction. Second, by explicitly including the state $x$ in the lifted representation (4), the model embeds a reconstruction constraint directly into the consistency loss. This coupling eliminates the need for a decoder, as prediction errors in the lifted space naturally reflect discrepancies in the original coordinates.
@@
Assumptions (A1)–(A2) are standard in online learning and naturally satisfied in our setting. The loss $\mathcal{L}_{t}$ is smooth in both $\theta$ and $K$, provided that the neural network $\tilde{\Phi}_{\theta}$ employs smooth activations. The bounded variation of the dynamic oracle (A2) reflects slowly evolving or piecewise-stationary dynamics, which commonly arise in practice. The key nonstandard assumption is (A3), which bounds the cumulative conformity thresholds $q_{t}$, and which follows from our use of conformal PI. More specifically, the function $h$ comes from the saturation function $r_{t}$ in the update rule (3), as designed in Angelopoulos2023ConformalPID. While we state (A3) as an assumption, it can be viewed as an empirical hypothesis: in regimes with stable distributions and smoothly adapting models, conformity thresholds decrease rapidly, leading to sublinear accumulation. This behavior is consistently observed across our experiments (see Fig. 2(a)), and deriving it remains an important direction for future work. For intuition, imagine a hypothetical scenario where (A3) does not hold, for instance where the cumulative sum grows linearly. This would imply that the scores stay roughly constant over time. From a conformal prediction standpoint, this means that predictive uncertainty does not shrink as more data arrives, indicating the model fails to learn a better representation. Such behavior may arise in non-autonomous systems under frequent abrupt changes, or when the model lacks the expressivity to learn the dynamics. While the former case is out of the scope of the present paper, the second case can be mitigated by augmenting the dimension of the lifted space.
Our proposed conformal online learning approach can be applied to more general online non-convex learning settings beyond Koopman framework. An important contribution to online non-convex learning is due to suggala2020online, who achieves optimal regret even under non-convex and adversarial losses by assuming access to an offline oracle. Subsequent work on dynamic regret explores variants to handle changing environments (xu2024online). Although these approaches specify how to update model parameters at each iteration, they do not consider when such updates should take place or to what extent. Our approach provides a statistically principled answer to these questions.
@@ 4.1 Illustration and validation
Setting. To validate the spectral accuracy of our online algorithm, we consider a benchmark system with analytically known Koopman eigenvalues and eigenfunctions (Brunton2016Control; Bruton2022ModernKoopman):
@@ 4.2 Comparison with online baselines
Datasets. We consider dynamical systems with a single attractor (Single Attractor(Brunton2016Control), two stable spirals and a saddle point (Duffing oscillator(Williams2015EDMD), a limit cycle (Van der Pol oscillator(sinha2019rRecursiveEDMD), and a chaotic regime with a strange attractor (Lorenz system(kostic2022rkhs). We complement them with three real-world datasets: the Electricity Transformer Dataset (ETD(zhou2021informer), the EEG Motor Movement/Imagery Dataset (2004EEG_1; 2000EEG_2) and a real turbulence dataset CASES-99 (CASES99). These datasets introduce additional challenges such as noise and potential distribution shifts. Full details are provided in Appendix B.1.
Baselines. We compare COLoKe against the online learning strategies listed in Table 1. While the table includes both online and batch-based methods, our evaluation focuses only on those compatible with fully streaming, one-sample-at-a-time updates, excluding methods that require access to data batches. For R-EDMD (sinha2019rRecursiveEDMD; sinha2023online), since the radial basis function dictionary requires the full data to estimate the centers (see Table 1), we use a polynomial dictionary. And the reconstruction matrix is estimated with the current buffer by regularized pseudo-inverse. This purely online variant is coined Online EDMD (OEDMD). For completeness, we also include a variant of COLoKe, referred to as OLoKe, in which the conformal update rule is dropped and replaced by a standard strategy: at each time step, the parameters are updated using a fixed number of gradient steps upon receiving a new sample. Implementation details are reported in Appendix B.2.
@@ 5 Conclusion
We have introduced COLoKe, a principled framework for online learning of Koopman-invariant representations, where conformal prediction is repurposed to adaptively trigger updates only when the model becomes temporally inconsistent. By moving beyond fixed-step updates, COLoKe achieves accurate and efficient learning of linear embeddings for nonlinear dynamics when data arrive sequentially. This work opens several promising directions for future research. On the theoretical front, our results highlight the need for a more comprehensive understanding of conformity-based online learning, with potential relevance well beyond Koopman operator estimation. The main limitation of our analysis is assumption (A3), whose validity remains open although being empirically supported; future work could aim to derive it from first principles for autonomous systems. On the methodological front, extending COLoKe to non-autonomous systems could further broaden its applicability and help unlock practical Koopman learning in real-time, resource-constrained, and dynamically evolving environments.

### v2

@@ 1 Introduction
Understanding the evolution of complex systems over time is essential in numerous disciplines, including robotics (Bruder et al., 2021), finance (Mann and Kutz, 2016), physics (Kaptanoglu et al., 2020), chemistry (Klus et al., 2020) and neuroscience  (Chrétien et al., 2023). A particularly powerful framework for this analysis is provided by operator-theoretic approaches, which recast nonlinear dynamics into linear evolution in function space. Among these, the Koopman operator (Budišić et al., 2012; Mauroy et al., 2020) plays a central role: it describes the progression of measurement functions (or observables) defined over the state space, yielding an infinite-dimensional linear representation of nonlinear systems. Its spectral decomposition offers a principled way to characterize the long-term behavior and underlying structure of the dynamics (Brunton et al., 2022).
@@
Although the Koopman operator is infinite-dimensional, several numerical methods have been developed to approximate it in a finite-dimensional setting. Notable among these are Dynamic Mode Decomposition (DMD) (Rowley et al., 2009) and its nonlinear extensions, such as Extended DMD (EDMD) (Williams et al., 2015) and its variants (see (Jin et al., 2024) and references therein). Recent work has also proposed truncated Signature features rooted in path theory (Chrétien et al., 2025). Another line of research consists of kernel learning formulations in reproducing kernel Hilbert spaces (Kostic et al., 2022; Hou et al., 2023). In addition, deep learning techniques have been leveraged to learn expressive Koopman-invariant representations, using neural networks and autoencoder architectures to estimate significant observables (Li et al., 2017; Wehmeyer and Noé, 2018; Lusch et al., 2018; Yeung et al., 2019; Otto and Rowley, 2019). For example, Xu et al. (2025b) addresses this by explicitly minimizing the spectral residual. The stochastic Koopman operator extends the classical Koopman framework to random dynamical systems, capturing the evolution of observables under stochastic influences (Črnjarić-Žic et al., 2020; Kostic et al., 2023; Xu et al., 2025a).
@@
Most existing methods, however, operate in the offline setting, assuming access to the entire dataset in advance. Yet, in many applications—such as online monitoring, adaptive control, or real-time forecasting—data arrive sequentially, and the system may evolve in a non-stationary fashion (Korda and Mezić, 2018). Several online methods have been proposed for this task. For instance, Online DMD (Zhang et al., 2019) and Online EDMD (Sinha et al., 2019, 2023) incrementally update Koopman operator estimates as new data arrives. However, these methods typically rely on linear observables or fixed dictionaries, limiting their expressiveness. More recent methods using neural networks (e.g. (Liang et al., 2022; Hao et al., 2024)) lack principled learning strategies and often rely on retraining the model using a fixed number of steps, regardless of whether the update is necessary. These limitations highlight the need for online learning strategies that are not only memory-efficient but also adaptive in order to update models only when required by the incoming data.
@@
We address this need by introducing Conformal Online Learning of Koopman embeddings (COLoKe), whose high-level principle is sketched in Figure 1. Our approach combines deep Koopman representation learning with a novel repurposing of conformal prediction principles to decide when to adapt the model as data arrive in a streaming fashion. Updates are triggered only when necessary, reducing both computational burden and overfitting. This contributes to the growing body of work on online Koopman learning (Sinha et al., 2023; Loya and Tallapragada, 2024; Mazouchi et al., 2023). Our method differs by enabling adaptive embeddings, built-in reconstruction, and real-time updates while remaining memory-efficient (see Table 1). To the best of our knowledge, COLoKe is the first principled approach for online learning driven by conformal-based updates.
@@
No need for history Online update Adaptative embedding Built-in reconstruction
ODMD (Zhang et al., 2019)
R-EDMD (Sinha et al., 2019, 2023)
DKLT (Hao et al., 2024) batch only
BatchOnline(Mazouchi et al., 2023) batch only
@@ 2.1 Data-driven learning of the Koopman operator
which describes the evolution of observables along trajectories of the system. Importantly, $\mathcal{K}$ is a linear operator, even when $T$ is nonlinear, making it a powerful tool for the spectral analysis of nonlinear dynamics (Brunton et al., 2022). In particular, if $\varphi\in\mathcal{F}$ is an eigenfunction of $\mathcal{K}$ with eigenvalue $\lambda\in\mathbb{C}$, i.e., $\mathcal{K}\varphi=\lambda\varphi$, then along a trajectory $(x_{t})$, the observable evolves linearly: $\varphi(x_{t})=\lambda^{t}\varphi(x_{0})$. It follows that, when a set of eigenfunctions $\{\varphi_{1},\dots,\varphi_{L}\}$ defines an injective embedding of the state space, the system can be linearized via the coordinate transformation $x\mapsto\left(\varphi_{1}(x),\dots,\varphi_{L}(x)\right)$. In this lifted space, the nonlinear dynamics evolve linearly, providing a compelling framework for Koopman-based analysis and control (Korda and Mezić, 2018; Mauroy et al., 2020).
@@
where $\Phi(x)=[f_{1}(x),\dots,f_{m}(x)]^{\top}$ denotes the lifted representation of the state. This insight motivates the search for low-dimensional Koopman-invariant subspaces that admit such linear representations. Methods like EDMD (Williams et al., 2015) approximate this setting by fixing a dictionary $\{f_{i}\}_{i=1}^{m}$, but their performance is limited by the expressiveness and suitability of the chosen observables. To overcome this limitation, recent approaches (Takeishi et al., 2017; Lusch et al., 2018; Yeung et al., 2019; Otto and Rowley, 2019) propose to learn both the feature map $\Phi$ and the linear operator $K$ jointly using neural networks, leading to deep Koopman embeddings.
@@ 2.2 Conformal prediction for online data
Conformal prediction (Vovk et al., 1999, 2005; Romano et al., 2019; Angelopoulos and Bates, 2021) is a distribution-free framework for uncertainty quantification that constructs valid prediction sets with finite-sample guarantees.
Given past input–output pairs $\{(x_{i},y_{i})\}_{i=1}^{t-1}$, a model produces a prediction $\hat{y}_{t}$ for a new input $x_{t}$, and assigns a conformity score $s(x_{t},y)$ to each candidate output $y$. The conformal prediction set is defined as $C_{t}=\left\{y\in\mathcal{Y}\;\middle|\;s(x_{t},y)\leq q_{t}\right\}$, where $q_{t}$ is a quantile calibrated to ensure $\mathbb{P}(y_{t}\in C_{t})\geq 1-\alpha$. The set $C_{t}$ is thus interpreted as a set of plausible outputs: it contains all candidate values of $y$ that are deemed sufficiently "conformal" (i.e., not too surprising) with respect to the current model and past observations. While this framework is powerful and requires no distributional assumptions beyond exchangeability, it is not directly applicable to time series or online learning, where data are typically non-exchangeable. In such settings, standard conformal methods may yield miscalibrated or overly conservative intervals. Addressing this challenge has motivated the development of adaptive conformal approaches that can track distribution shifts over time (Gibbs and Candes, 2021; Xu et al., 2021; Angelopoulos et al., 2023). In particular, Conformal PID Control was recently introduced in Angelopoulos et al. (2023) as a dynamically calibrated version of conformal prediction. Here, “PID” refers to the use of Proportional, Integral, and (optionally) Derivative feedback terms—standard components in control theory—used to adaptively adjust the prediction threshold. Rather than fixing the quantile threshold $q_{t}$ in advance, the method updates it online in response to conformity violations. Among its variants, we consider the conformal PI control scheme: after observing whether $y_{t}\in C_{t}$, a binary error signal $e_{t}=\mathbf{1}\{y_{t}\notin C_{t}\}$ is computed, and the threshold is adjusted via:
@@
where $\gamma>0$ is a learning rate and $r_{t}$ is a nonlinear saturation function (Angelopoulos et al., 2023) acting as an integral correction term. This PI-style control loop allows the prediction region to adaptively expand or contract to maintain the desired coverage. Under mild regularity conditions, it ensures that the empirical coverage converges, i.e., $\frac{1}{T}\sum_{t=1}^{T}e_{t}\to\alpha$ as $T\to\infty$.
@@ 3.1 From online learning…
At each time step $t$, we denote by $\Phi_{\theta_{t}}:\mathcal{X}\to\mathbb{C}^{m}$ the current feature map and by $K_{t}\in\mathbb{C}^{m\times m}$ the corresponding finite-dimensional Koopman operator. Following (Li et al., 2017), we also advocate to enforce interpretability and preserve part of the original state, by designing the feature map to include both the identity and a learnable nonlinear component, i.e.,
@@
This formulation is motivated by two design principles. First, as shown by Otto and Rowley (2019), multi-step prediction promotes the identification of persistent spectral modes and approximate Koopman eigenfunctions, thereby improving long-term prediction. Second, by explicitly including the state $x$ in the lifted representation (4), the model embeds a reconstruction constraint directly into the consistency loss. This coupling eliminates the need for a decoder, as prediction errors in the lifted space naturally reflect discrepancies in the original coordinates.
@@
Assumptions (A1)–(A2) are standard in online learning and naturally satisfied in our setting. The loss $\mathcal{L}_{t}$ is smooth in both $\theta$ and $K$, provided that the neural network $\tilde{\Phi}_{\theta}$ employs smooth activations. The bounded variation of the dynamic oracle (A2) reflects slowly evolving or piecewise-stationary dynamics, which commonly arise in practice. The key nonstandard assumption is (A3), which bounds the cumulative conformity thresholds $q_{t}$, and which follows from our use of conformal PI. More specifically, the function $h$ comes from the saturation function $r_{t}$ in the update rule (3), as designed in Angelopoulos et al. (2023). While we state (A3) as an assumption, it can be viewed as an empirical hypothesis: in regimes with stable distributions and smoothly adapting models, conformity thresholds decrease rapidly, leading to sublinear accumulation. This behavior is consistently observed across our experiments (see Fig. 2(a)), and deriving it remains an important direction for future work. For intuition, imagine a hypothetical scenario where (A3) does not hold, for instance where the cumulative sum grows linearly. This would imply that the scores stay roughly constant over time. From a conformal prediction standpoint, this means that predictive uncertainty does not shrink as more data arrives, indicating the model fails to learn a better representation. Such behavior may arise in non-autonomous systems under frequent abrupt changes, or when the model lacks the expressivity to learn the dynamics. While the former case is out of the scope of the present paper, the second case can be mitigated by augmenting the dimension of the lifted space.
Our proposed conformal online learning approach can be applied to more general online non-convex learning settings beyond Koopman framework. An important contribution to online non-convex learning is due to Suggala and Netrapalli (2020), who achieves optimal regret even under non-convex and adversarial losses by assuming access to an offline oracle. Subsequent work on dynamic regret explores variants to handle changing environments (Xu and Zhang, 2024). Although these approaches specify how to update model parameters at each iteration, they do not consider when such updates should take place or to what extent. Our approach provides a statistically principled answer to these questions.
@@ 4.1 Illustration and validation
Setting. To validate the spectral accuracy of our online algorithm, we consider a benchmark system with analytically known Koopman eigenvalues and eigenfunctions (Brunton et al., 2016, 2022):
@@ 4.2 Comparison with online baselines
Datasets. We consider dynamical systems with a single attractor (Single Attractor(Brunton et al., 2016), two stable spirals and a saddle point (Duffing oscillator(Williams et al., 2015), a limit cycle (Van der Pol oscillator(Sinha et al., 2019), and a chaotic regime with a strange attractor (Lorenz system(Kostic et al., 2022). We complement them with three real-world datasets: the Electricity Transformer Dataset (ETD(Zhou et al., 2021), the EEG Motor Movement/Imagery Dataset (Schalk et al., 2004; Goldberger et al., 2000) and a real turbulence dataset CASES-99 (Earth Observing Laboratory, 1999). These datasets introduce additional challenges such as noise and potential distribution shifts. Full details are provided in Appendix B.1.
Baselines. We compare COLoKe against the online learning strategies listed in Table 1. While the table includes both online and batch-based methods, our evaluation focuses only on those compatible with fully streaming, one-sample-at-a-time updates, excluding methods that require access to data batches. For R-EDMD (Sinha et al., 2019, 2023), since the radial basis function dictionary requires the full data to estimate the centers (see Table 1), we use a polynomial dictionary. And the reconstruction matrix is estimated with the current buffer by regularized pseudo-inverse. This purely online variant is coined Online EDMD (OEDMD). For completeness, we also include a variant of COLoKe, referred to as OLoKe, in which the conformal update rule is dropped and replaced by a standard strategy: at each time step, the parameters are updated using a fixed number of gradient steps upon receiving a new sample. Implementation details are reported in Appendix B.2.
@@ 5 Conclusion
We have introduced COLoKe, a principled framework for online learning of Koopman-invariant representations, where conformal prediction is repurposed to adaptively trigger updates only when the model becomes temporally inconsistent. By moving beyond fixed-step updates, COLoKe achieves accurate and efficient learning of linear embeddings for nonlinear dynamics when data arrive sequentially. This work opens several promising directions for future research. On the theoretical front, our results highlight the need for a more comprehensive understanding of conformity-based online learning, with potential relevance well beyond Koopman operator estimation. The main limitation of our analysis is assumption (A3), whose validity remains open although being empirically supported; future work could aim to derive it from first principles for autonomous systems. On the methodological front, extending COLoKe to non-autonomous systems (Gao et al., 2025) could further broaden its applicability and help unlock practical Koopman learning in real-time, resource-constrained, and dynamically evolving environments.